OpenStudy (anonymous):

w-16=6 and 1 over five

OpenStudy (anonymous):

What are you supposed to find?

OpenStudy (anonymous):

w

OpenStudy (anonymous):

well the first one you need to isolate w. so add 16 to both sides to get w = 16+6, leaving w = 22

OpenStudy (anonymous):

Not sure where w plays into the 2nd part

OpenStudy (anonymous):

6 11/30

OpenStudy (anonymous):

What are you supposed to do with that?

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

I'm not sure what exactly your problem means then

OpenStudy (anonymous):

but w-16=6 will leave you with w = 22

OpenStudy (amistre64):

w - 16 = 6 and 1/5 ...its a mixed numeral

OpenStudy (anonymous):

So "and" means multiplied by?

OpenStudy (anonymous):

or added to?

OpenStudy (amistre64):

31/5 is the improper fractions of 6 and 1/5

OpenStudy (amistre64):

integer 6 and fraction 1/5

OpenStudy (amistre64):

6.2 as a decimal :)

OpenStudy (anonymous):

Yeah that makes sense now. I wish people would be a bit more clear about their problems

OpenStudy (amistre64):

it stumped me for a second too ;)

OpenStudy (anonymous):

well you should know how to read it it was written in simplest form. dont get mad because you didnt know the answer

OpenStudy (anonymous):

i wrote it the way it was given